The convex lens of a copy machine has a focal length of 25cm. A letter to be copied is placed 40 cm from the lens. How far from the lens is the copy paper located? How much larger will the copy be?
1
Expert's answer
2011-06-25T07:48:54-0400
1/S1 +1/S2 =1/f

1/S2 = -1/S1 +1/f = (S1-f)/(S1 f)
S2 = (S1 f)/(S1-f) = (40*25)/(40-25) = 1000/15 = 67cm.

Magnification is
M = -S2/S1 = -67/40 = -1.675.
Notice the sign convention here shows that, if M is negative, as it is for real images, the image is upside-down with respect to the object. For virtual images, M is positive and the image is oriented upright.